OpenStudy (simplymarie_x):

I need more opinions because people keep telling me it's not dimensional analysis but others are telling me it is. PLEASE HELP. Use dimensional analysis to find the number of cm/min in 5.8 m/hr. Show your work and be sure to calculate the answer in the correct number of significant figures.

OpenStudy (irishboy123):

it's not really dimensional analysis. dimensional analysis helps ypu connect variables by looking at the units they are measured in. so you could take Newtons and mass and guess some sort of relationship with acceleration that had to be linear because of the dimensions they are measured in. the constant clearly would be up for grabs. here you are just doing a conversion between units. however you can still say that : \(\dfrac{cm}{min} = \dfrac{m/100}{hr/60}\) or \(\dfrac{m}{hr} = \dfrac{cm*100}{min*60}\) or something like that .... but that's not dimensional analysis as i understand the term.......:-(